Suppose that a persons yearly income is $80,000. Also suppose that this persons money demand function is given by M-SY050-
Answer #1

Yearly income = $80,000

Interest rate = 15% or 0.15

Md = $Y(0.50 - i)

Md = $80,000(0.50 - 0.15) = $80,000 * 0.35 = $28,000

Now, yearly income decreases by 25%.

new yearly income = $80,000 - ($80,000 * 0.25) = $80,000 - $20,000 = $60,000

New money demand = $Y(0.50 - i) = $60,00(0.50 - 0.15) = $60,000 * 0.35 = $21,000

Percentage change in money demand = [($28,000 - $21,000)/$28,000] * 100 = 25%

So,

Suppose that the interest rate is 15%.

The percentage change in this person's demand for money if the yearly income falls by 25% is 25%.

Yearly income = $80,000

Interest rate = 10% or 0.10

Md = $Y(0.50 - i)

Md = $80,000(0.50 - 0.10) = $80,000 * 0.40 = $32,000

Now, yearly income decreases by 25%.

new yearly income = $80,000 - ($80,000 * 0.25) = $80,000 - $20,000 = $60,000

New money demand = $Y(0.50 - i) = $60,00(0.50 - 0.10) = $60,000 * 0.40 = $24,000

Percentage change in money demand = [($32,000 - $24,000)/$32,000] * 100 = 25%

So,

Suppose that the interest rate is 10%.

The percentage change in this person's demand for money if the yearly income falls by 25% is 25%.

The above calculations shows that any decrease (or increase) in income leads to a proportional increase in money demand regardless of the interest rate.

Hence, the correct answer is the option (C).